What is the coefficient of x^5 in y = Sinh(X)


1
Expert's answer
2022-01-31T18:06:05-0500

Solution:

We know that,

"\sinh x=x+\dfrac{x^3}{3!}+\dfrac{x^5}{5!}+\dfrac{x^7}{7!}+\dfrac{x^9}{9!}+..."

Then, coefficient of "x^5=\dfrac{1}{5!}=\dfrac{1}{120}"